Train A has a speed 15 miles per hour greater than that of train B. If train A travels 270miles in the same times train B travels
240 miles, what are the speeds of the two trains?
Train A was traveling at ( ) miles an hour where train B was traveling at ( ) miles per hour???

Answer by rothauserc(4718)

rate * time = distance
:
Let the rate of train B be r, then rate of train A is r+15
:
(1) (r+15) * t = 270
:
(2) r * t = 240
:
There are two equations in two unknowns
:
solve equation 2 for t
:
t = 240/r
:
substitute for t in equation 1
:
(r+15) * 240/r = 270
:
240r +3600 = 270r
:
30r = 3600
:
r = 120
:
Train B travels at 120 miles per hour
:
Train A travels at 120 +15 = 135 miles per hour
:
*****************************************************************************
Train A was traveling at 135 miles an hour where train B was traveling at 120 miles per hour
*****************************************************************************
: